# Self-supervised Learning of Event-guided Video Frame Interpolation for Rolling Shutter Frames

### v1

@@ Abstract
This paper makes the first attempt to tackle the challenging task of recovering arbitrary frame rate latent global shutter (GS) frames from two consecutive rolling shutter (RS) frames, guided by the novel event camera data. Although events possess high temporal resolution, beneficial for video frame interpolation (VFI), a hurdle in tackling this task is the lack of paired GS frames. Another challenge is that RS frames are susceptible to distortion when capturing moving objects. To this end, we propose a novel self-supervised framework that leverages events to guide RS frame correction and VFI in a unified framework. Our key idea is to estimate the displacement field (DF) —non-linear dense 3D spatiotemporal information of all pixels during the exposure time, allowing for the reciprocal reconstruction between RS and GS frames as well as arbitrary frame rate VFI. Specifically, the displacement field estimation (DFE) module is proposed to estimate the spatiotemporal motion from events to correct the RS distortion and interpolate the GS frames in one step. We then combine the input RS frames and DF to learn a mapping for RS-to-GS frame interpolation. However, as the mapping is highly under-constrained, we couple it with an inverse mapping (i.e., GS-to-RS) and RS frame warping (i.e., RS-to-RS) for self-supervision. As there is a lack of labeled datasets for evaluation, we generate two synthetic datasets and collect a real-world dataset to train and test our method. Experimental results show that our method yields comparable or better performance with prior supervised methods.
Figure 1: We propose a novel VFI method that enables the recovery of in-between GS frames (d) with an arbitrary frame rate from two consecutive RS frames (a and b), guided by the novel event camera data (c) in a self-supervised manner. Here, 0 - 0.5 and 0.5 - 1 are the normalized exposure time. Importantly, our method can achieve a significant temporal resolution up to 32 $\times$ frame rate (e).
@@
1 Introduction
CMOS cameras are dominantly applied to consumer-grade devices, e.g., mobile phones, due to their low cost and power consumption [8, 3, 16, 36]. These sensors, however, struggle with the rolling shutter (RS) effect, leading to unexpected distortion problems, e.g., skew, when capturing fast-moving objects. By contrast, global shutter (GS) sensors synchronize the exposure period of all pixels; therefore, they do not suffer from edge distortion and region occlusion [26, 8, 5, 14, 23]. With the growing demand for high frame rate and high-quality videos for mobile devices, deep learning-based video frame interpolation (VFI) has attracted increasing attention in the research community [18, 41]. Despite the success of these VFI methods  [25, 18, 27, 42, 28, 2, 32, 30], they all follow the GS mechanism. Therefore, they often fail to recover latent GS frames from two consecutive RS frames, captured by the RS sensors, especially, the high-speed or dynamic motion scenes [38].To tackle this problem, many learning-based RS correction methods [24, 34, 47] have been proposed to obtain GS frames by removing the RS effect based on the assumption of linear motion. However, they can not synthesize nonexistent in-between GS frames. Therefore, it is desirable to generate in-between GS frames from two consecutive RS frames as it can benefit both VFI and RS correction.Accordingly, some learning-based methods have been proposed. RSSR [5] and CVR [8] are two representative methods that estimate the linear motion to recover faithful in-between GS frames from two consecutive RS frames. However, they can not handle non-linear motion, which often occurs in dynamic scenes with fast motion.
@@ 1 Introduction
In this paper, we make the first attempt to leverage the high temporal resolution of events to guide the recovery of in-between GS frames from two consecutive RS frames. However, tackling this novel problem is non-trivial because 1) there do not exist GS-events-RS triplet datasets for VFI, and 2) RS frames are suspectable to edge distortion and region occlusion in dynamic scenes with fast motion. To this end, we propose a novel self-supervised learning (SSL) framework that leverages events to guide RS correction and VFI in a unified framework. Overall, our method enables the recovery of GS frames with any arbitrary frame rate, e.g., 32$\times$, from two consecutive RS frames, guided by events, as depicted in Fig. Self-supervised Learning of Event-guided Video Frame Interpolation for Rolling Shutter Frames. The proposed SSL framework is shown in Fig. 2. The key idea of our method is to 1) estimate the 3D displacement field (DF), which includes dense spatiotemporal non-linear motion information of all pixels during the exposure time, and 2) combine the RS frames and DF for the reciprocal reconstruction (or mapping)to impose self-supervision.
@@
Specifically, we first propose the displacement field estimation module to estimate the spatiotemporal motion information from events directly (Sec. 4.1).We split events into moments, each of which includes a fixed amount of voxel grid [33, 11].We estimate the optical flow [12] between consecutive event moments. This way, we can obtain the non-linear 3D motion information—DF, during the exposure time for GS frames.Benefiting from the high temporal resolution of events, 3D DF contains dense motion information for RS correction and GS frames interpolation in one step.Based on DF, we propose a latent GS frames generation module to learn the RS-to-GS mapping for GS frame interpolation (Sec. 4.2). We generate a series of GS frames at arbitrary frame rate in exposure time from RS frames and 3D DF. As the ground truth GS frames are not available, the mapping is highly under-constrained. Thus, we couple it with a reciprocal reconstruction module to 1) reconstruct RS frames based on the generated GS frames and fully exploit the constraints inherent in RS frames, and 2) achieve RS-to-RS warping based on the DF for self-supervision (Sec. 4.3).
@@
Figure 2: Overview of our proposed framework, which consists of three parts, the displacement field estimation module, latent GS frames generation module, and the reciprocal reconstruction module. See contexts for more details.
@@
Due to the lack of RS-events-GS triplet datasets, we generate two simulated datasets for qualitative and quantitative evaluation. Moreover, we collect a real-world dataset with RS frames and aligned events for training and qualitative assessment of our SSL method.Experimental results show that our method yields 1) comparable performance with supervised frame-based RS-to-GS VFI method [8] and 2) better performance than the supervised event-guided RS correction [46] + unsupervised event-guided VFI method—TimeReplayer [15]. Note that with RS frames and events as inputs, our method achieves much higher VFI performance than that of TimeLens (only for GS frames) [38].In summary, the contributions of this paper are four-fold: (I) we make the first attempt to address GS frame VFI from RS frames, guided by events. (II) We propose an SSL framework that can generate arbitrary frame rate GS frames from RS frames and events. (III) We introduce a DF estimation module and a reciprocal reconstruction module to impose self-supervision. (IV) We introduce two synthetic and one real-world datasets for the training and evaluation.
Methods Event Based Self Supervised Rolling Correction Frame Interpolation
EvUnroll [46]
TimeReplayer [15]
TimeLens [38]
RSSR  [5]
CVR  [8]
Ours
Table 1: Comparison of our method with prior-arts.
2 Related Works
@@
Event-guided VFI: Video frame interpolation (VFI) is a fundamental task converting low frame rate videos to high frame rate videos in video enhancement [25, 4, 31]. During generating intermediate frames between consecutive frames, frame-based methods [25, 18, 27, 42, 28, 2, 32, 30] predict motion explicitly or implicitly. However, these methods degrade greatly in scenes of non-linear motion since accurately modeling motion from the sparse set of frames is ill-posed [43]. Unlike the RGB camera, event camera enjoys many advantages, such as high temporal resolution, which captures the brightness changes in a time interval [38]. Previous works have demonstrated its potential for VFI, and they can be roughly categorized into two types: supervised and unsupervised methods.The supervised methods can also be divided into two parts: synthesis-based and wrapping-based methods. For example, EFI [29], is a synthesis-based method that generates intermediate frames from events and images. $A^{2}$ [41] explores the relation between event counts and motion, and designs event-driven optical flow masks for warping intermediate frames. Time Lens [38] and Time Lens++ [37] combine synthesis methods with warping-based methods to boost the VFI performance.TimeReplayer [15] is an unsupervised approach that applies a loss between the input frames and reconstructed input frames—warped from interpolated frames. All these methods are designed for the GS frames without considering the distortion caused by RS. However, most commercial cameras record frames with the RS mechanism, thus impeding their applications in real-world scenarios.
@@
RS correction:Recently, some learning-based methods [24, 6, 7, 1, 8] have been proposed to achieve the RS correction. As the motion information between frames is unknown, these methods rely on the prior assumption of linear motion to predict the intermediate GS frames, which degrade greatly in scenes of non-linear motion.Zhou et al. [46] explored the spatiotemporal information of event cameras to boost the performance of RS correction with the non-linear motion for the target time. However, they focus on RS correction and deblurring without considering RS-to-GS frame interpolation.
@@
VFI with RS frames: RSSR [5] proposes the first work to recover a random frame rate GS frame from RS frames, while the results suffer from unwanted holes and black edges, caused by occlusion between the RS and GS frames.CVR [8] then proposes a context-aware GS frame interpolation framework to alleviate the occlusion problem and reduce artifacts. However, these methods are frame-based and only focus on linear motion. It is demanding to consider non-linear motion for real-world applications. In addition, these methods need the paired RS-GS dataset for training, which is difficult to collect. Therefore, they only conduct experiments on the synthetic dataset.We make the first attempt to leverage events to guide RS correction and VFI in one step, as shown in Tab. 1. Accordingly, we propose a self-supervised approach that generates the arbitrary frame rate GS frames between consecutive RS frames.
3 Preliminaries
@@
Rolling Shutter is an exposure mechanism widely deployed in commercial cameras, which determines the VFI problem definition.First, we denote $\bm{L}$ as the intensity of an ideal video sequence in the period of $[t_{s},t_{e}]$ containing continuous GS frames.A RS frame $I_{r}$ can be regarded as a special composition of a series of GS frames. We denote the $t^{r}_{s}$ and $t^{r}_{e}$ as the start and end of the exposure time of RS frame $I_{r}$, respectively. The interval of exposure time between every two adjacent lines is $\Delta t=(t^{r}_{s}-t^{r}_{e})/(H-1)$, where $H$ is the height of video frames. Therefore, the exposure time of row $h$ of the RS frame can be recorded as $t_{h}=t^{r}_{s}+h\times\Delta t$. Then the raw $h$ of RS frame $I_{r}$ can be formulated as Eq. 1, where $I_{t_{h}}[h]$ refers to the row $h$ of the GS frame captured at $t_{h}$.
@@
4 Proposed Framework
Overview:The overall framework of the proposed approach is depicted in Fig. 2, which can be divided into three parts: (I) Displacement Field Estimation, (II) Latent GS frames Generation, and (III) Reciprocal Reconstruction. As the oversimplified linear motion model fails in complex non-linear motions, we first introduce DF which contains non-linear motion during the exposure time and bridge the gap between RS frames and GS frames (Sec. 4.1), followed by describing how to generate latent GS frames by RS frames and DF (Sec. 4.2).Since the mapping from RS to GS is highly under-constrained, we couple it with the inverse mapping (GS-to-RS) and RS frame warping (RS-to-RS) for self-supervision (Sec. 4.3). The inputs of our framework are two consecutive RS frames($I_{r_{0}}$, $I_{r_{1}}$) and their corresponding events, while the outputs are continuous latent global shutter (GS) frames at arbitrary frame rate, as shown in Fig. 3.
4.1 Displacement Field Estimation (DFE)
@@
To address this issue, we leverage the high-temporal resolution of events to estimate the optical flow within $[t_{0},t_{1}]$. We divide the events into $T+1$ time bins, and each time bin is further divided into $N$ voxel grids [33, 39]. Based on this, we can get the event representation with a dimension of $(T+1)\times N\times H\times W$. We denote the event-based optical flow estimation function as $f_{flow}$, and the dimension of the estimated optical flow set is $2\times T\times H\times W$. In practice, we estimate a set of optical flows as the $\mathcal{D}$ by E-RAFT  [12], as it has shown promising results in handling challenging motion scenes. In this way, we can fit the nonlinear motion with the linear motions of $T$ segments.If we know the location of a pixel at time $t_{0}$ and there is a considerable time gap between $t_{0}$ and $t_{n}$, we can obtain the location of this point at time $t_{n}$ by adding up its displacement in DF over the time interval, as Eq. 2.
@@
$\mathcal{L}_{field}=\frac{1}{T}\sum_{i=1}^{T}\left((\nabla_{x}\mathcal{D}[i])^{2}+(\nabla_{y}\mathcal{D}[i])^{2}\right)$ (3)
@@
Figure 4: An illustration of masks of RS-to-GS and RS-to-RS for the displacement field. Blue represents a positive value, indicating forward movement. Red represents a negative value, indicating backward movement. Each item in the mask uses uniform sample $S_{h}\times S_{t}$ points to estimate its weight. (a) shows more detail. Two-time points control the rolling shutter frame, the start exposure time $t_{rs}$ and the end exposure time $t_{re}$. The global shutter frame is controlled by the exposure time $t_{g}$. (d) (e) (f) shows the projection of different masks on the $H\times T$ dimension. For example, (d) shows the transition from rolling shutter frame (start exposure time 0 and end exposure time 0.7) to global shutter frame (exposure time 0.4).
@@
Mask Design for RS GS transformation We designed the mask based on the exposure mechanism of the rolling shutter and global shutter to describe the transition between any two frames. The real exposure model of the rolling shutter frame and events are shown in Fig.4. The red solid line represents the start time of the frame exposure, and the red dashed line represents the end time of the frame exposure. It can be easily found that the exposure time is much shorter than the rolling shutter time. Long exposure time can lead to blurring, which complicates the research question of frame interpolation, and we do not consider the long exposure in the paper. Therefore, we approximate the rolling shutter frame and the global shutter frame as planes. Obviously, the plane of the global shutter can be described as $t=t_{g}$, where $t_{g}$ is the global shutter exposure timestamp. The plane of the rolling shutter is parallel to the W-axis and passes through two points $(h,t_{rs})$ and $(0,t_{re})$ at the same time, where $h$ is the height of frames, and $t_{rs}$ and $t_{re}$ are begin time and end time of rolling shutter frame exposure.
Given two planes, we use uniform sampling to calculate the weights of each bin in each mask. For each sampling point, we calculate whether it is on the left or right of the plane by computational geometry. Fig.4 shows the projection of more masks in the $H\times T$ dimension.
@@
Figure 5: The illustration of latent GS generation from RS frames $I_{r_{0}}$, $I_{r_{1}}$ and displacement field $\mathcal{D}$. The dash boxes indicate the output.
4.2 Latent Global Shutter Frame Generation

### v2

@@ Abstract
Index Terms:
Event camera, frame interpolation, rolling shutter, global shutter, bandwidth-efficient vision.
publicationid: pubid:
@@
I Introduction
The burgeoning interest in virtual reality (VR) has recently been extended to most consumer cameras, e.g.unmanned aerial vehicles (UAVs), offering users an unprecedented and immersive flight experience [5, 6, 7, 8]. High-frame-rate (i.e., slow-motion) videos free from distortions are pivotal for achieving this level of immersion [9, 10, 11, 12, 13]. However, attaining such an ideal video is constrained by inherent hardware limitations [14, 15, 16]. Specifically, the rolling shutter (RS) sensors, which are commonly used in most consumer cameras, introduce distortions, e.g., skew and the jelly effect, particularly during rapid movements [17, 18, 19, 20]. While the global shutter (GS) sensors circumvent RS distortion, they come at a cost that is more than ten times higher than RS sensors and also consume significant power [11]. Additionally, constraints on bandwidth [21, 22] and computational speed [23, 10, 22, 24, 25] also should be considered for VR video streaming. Against this backdrop, we aim to enhance video frame rate and correct RS distortion by introducing a framework characterized by fast inference speed and low bandwidth.
In light of these hardware constraints, there has been a growing demand for software-based solutions to achieve high-quality, high-frame-rate videos. This has propelled deep learning-based video frame interpolation (VFI) methods to the forefront of research [26, 27]. However, it’s worth noting that despite the success of these VFI methods [28, 26, 29, 30, 31, 32, 33, 34], they predominantly operate under the GS mechanism, which doesn’t directly address the RS distortions commonly found in UAV-captured videos, especially the high-speed or dynamic motion scenes [3, 35]. To tackle this problem, learning-based RS correction methods [36, 37, 38] have been proposed to obtain GS frames by removing the RS effect based on the assumption of linear motion. However, they can not synthesize nonexistent in-between GS frames. Therefore, it is desirable to generate in-between GS frames from two consecutive RS frames, as it can benefit both VFI and RS correction. Accordingly, some learning-based methods have been proposed. RSSR [39] and CVR [17] are two representative methods that estimate the linear motion to recover faithful in-between GS frames from two consecutive RS frames. However, they can not handle non-linear motion, which often occurs in dynamic scenes with fast motion.
[…]
@@ I Introduction
In this paper, we make the first attempt to leverage the high temporal resolution of event cameras to guide the recovery of in-between GS frames from two consecutive RS frames based on a multi-sensor-equipped UAV. However, tackling this novel problem is non-trivial because 1) there are no GS-events-RS triplet datasets for VFI, and 2) RS frames are suspectable to edge distortion and region occlusion in dynamic scenes with fast motion. To this end, we propose a novel self-supervised learning (SSL) framework that leverages events to guide RS correction and VFI in a unified framework. Overall, our method enables the recovery of GS frames with any arbitrary frame rate, e.g., 32$\times$, from two consecutive RS frames, guided by events, as depicted in Fig. 1. The proposed SSL framework is shown in Fig. 2. The key idea of our method is to 1) estimate the 3D displacement field (DF), which includes dense spatiotemporal non-linear motion information of all pixels during the exposure time, and 2) combine the RS frames and DF for the reciprocal reconstruction (or mapping) to impose self-supervision.
@@
Specifically, we first propose the displacement field estimation module to estimate the spatiotemporal motion information from events directly (Sec. IV-A). We split events into moments, each of which includes a fixed amount of voxel grid [52, 53, 54]. We estimate the optical flow [55] between consecutive event moments. This way, we can obtain the non-linear 3D motion information—DF, during the exposure time for GS frames. Benefiting from the high temporal resolution of events, 3D DF contains dense motion information for RS correction and GS frames interpolation in one step. Based on DF, we propose a latent GS frames generation module to learn the RS-to-GS mapping for GS frame interpolation (Sec. IV-B). We generate a series of GS frames at arbitrary frame rate in exposure time from RS frames and 3D DF. As the ground truth GS frames are not available, the mapping is highly under-constrained. Thus, we couple it with a reciprocal reconstruction module to 1) reconstruct RS frames based on the generated GS frames and fully exploit the constraints inherent in RS frames, and 2) achieve RS-to-RS warping based on the DF for self-supervision (Sec. IV-C).
@@
Figure 2: Overview of our proposed framework, which consists of three parts, (I) the displacement field estimation module, (II) latent GS frames generation module, and (III) the reciprocal reconstruction module. Additional context and explanations are provided in the main text.
@@
Due to the lack of RS-events-GS triplet datasets, we generate two simulated datasets for qualitative and quantitative evaluation. We propose the first real-world dataset with RS frames and aligned events for training and evaluation of our framework. Also, we collect a dataset using a UAV (equipped with RGB and event cameras) to validate our framework’s effectiveness. We conduct objective and subjective evaluations for our framework, focusing on key metrics such as inference speed, bandwidth efficiency, generalization performance. The evaluation results show that our method yields 1) comparable performance with supervised frame-based RS-to-GS VFI method [17] and 2) better performance than the supervised event-guided RS correction [4] + unsupervised event-guided VFI method—TimeReplayer [46]. Note that with RS frames and events as inputs, our method achieves much higher VFI performance than that of TimeLens (only for GS frames) [3]. In practice, the results underscore the remarkable prowess of our method in restoring slow-motion videos to their pristine quality, achieving this feat with a remarkable $94\%$ reduction in bandwidth usage and $16~{}ms$ per frame inference speed in the demanding scenario of 32×32$\times$32 × frame interpolation for UAV-VR application purpose.
In summary, the contributions of this paper are four-fold: (I) we propose the first self-supervised approach to recover GS slow-motion video from a multi-sensor-equipped UAV. (II) We introduce a DF estimation module and a reciprocal reconstruction module to impose self-supervision. (III) We propose the first real-world RS-event paired dataset for the training and evaluation. (IV) We conduct analysis of our approach, showcasing the benefits of our method with respect to bandwidth efficiency and inference speed. Collectively, our approach serves as a technical exploration, opening up possibilities for improving the UAV-VR video streaming experience by correcting RS distortion and enhancing frame rate.
II Related Works
The growing fascination with VR has expanded its horizons into the realm of UAVs, promising users an unparalleled and immersive flying experience [5, 6, 7, 56]. Achieving a heightened sense of immersion in this context heavily relies on the availability of high-frame-rate (i.e., slow-motion) videos devoid of distortions  [9, 10, 11, 12, 13]. Our primary focus is on harnessing events to guided RS frame correction and interpolation, leading us to categorize recent technical research into the following three facets:
@@ II Related Works
Event-guided VFI: Video frame interpolation (VFI) is a fundamental task converting low frame rate videos to high frame rate videos in video enhancement [28, 57, 58]. During generating intermediate frames between consecutive frames, frame-based methods [28, 26, 29, 30, 31, 32, 33, 34] predict motion explicitly or implicitly. However, these methods degrade greatly in scenes of non-linear motion since accurately modeling motion from the sparse set of frames is ill-posed [48]. Unlike the RGB camera, the event camera enjoys many advantages, such as high temporal resolution, which captures the brightness changes in a time interval [3]. Previous works have demonstrated its potential for VFI, and they can be roughly categorized into two types: supervised and unsupervised methods. The supervised methods can also be divided into two parts: synthesis-based and wrapping-based methods [47, 27]. Time Lens [3] and Time Lens++ [51] combine synthesis methods with warping-based methods to boost the VFI performance. TimeReplayer [46] is an unsupervised approach that applies a loss between the input frames and reconstructed input frames—warped from interpolated frames. All these methods are designed for the GS frames without considering the distortion caused by RS. However, most commercial cameras record frames with the RS mechanism, thus impeding their applications in real-world scenarios.
@@
RS correction: Recently, some learning-based methods have been proposed to achieve the RS correction [59, 36, 60, 61, 62, 17]. As the motion information between frames is unknown, these methods rely on the prior assumption of linear motion to predict the intermediate GS frames, which degrade greatly in scenes of non-linear motion. Zhou et al[4] explored the spatiotemporal information of event cameras to boost the performance of RS correction with the non-linear motion for the target time. However, they focus on RS correction and deblurring without considering RS-to-GS frame interpolation.
@@
VFI with RS frames: RSSR [39] proposes the first work to recover a random frame rate GS frame from RS frames, while the results suffer from unwanted holes and black edges, caused by occlusion between the RS and GS frames. CVR [17] then proposes a context-aware GS frame interpolation framework to alleviate the occlusion problem and reduce artifacts. However, these methods are frame-based and only focus on linear motion. It is demanding to consider non-linear motion for real-world applications. In addition, these methods need the paired RS-GS dataset for training, which is difficult to collect. Therefore, they only conduct experiments on the synthetic dataset. We make the first attempt to leverage events to guide RS correction and VFI in one step. Accordingly, we propose a self-supervised approach that generates the arbitrary frame rate GS frames between consecutive RS frames.
TABLE I: Definitions of Mathematical Symbols
Symbol Definition
$\bm{L}$ Ideal Continuous Global Shutter (GS) Video Intensity
$I_{r}$ Rolling Shutter (RS) Frame
$t^{r}_{s}$, $t^{r}_{e}$ RS Frame Exposure Start/End Times
$\Delta t$ RS Line Exposure Interval
$H$ Video Frame Height
$t_{h}$ RS Row $h$ Exposure Time
$E$ Event Stream
$e$ Single Event $(t,x,y,p)$
$C$ Event Triggering Threshold (Brightness)
$\bm{L}(t,x,y)$ Brightness at Pixel $(x,y)$ at Time $t$
$\Delta b$ Brightness Change
$\Phi$ Event Triggering Function
$p_{e}$ Brightness Change Polarity (+/-)
$tr(t)$ Point Trajectory $(x,y)$ during Exposure
$\Delta p_{0\to 1}$ Displacement from Time $t_{0}$ to $t_{1}$
$f_{flow}$ Event-based Optical Flow Estimation Function
$\mathbf{D}$ Displacement Field
$\mathbf{L}_{field}$ Displacement Field Loss
$\nabla$ Directional Gradient
$\mathbf{M}$ Weight Map for RS/GS Transformation
$f_{warp}$ Warping Function
$\mathbf{L}_{gs2rs}$ GS-to-RS Self-Supervision Loss
$\mathbf{L}_{rs2rs}$ RS-to-RS Self-Supervision Loss
$\mathbf{L}$ Total Loss (Weighted Sum)
III Preliminaries
@@
Rolling Shutter is an exposure mechanism widely deployed in commercial cameras, which determines the VFI problem definition. For clarity, the mathematical symbols used in this paper are defined in Tab. I. First, we denote $\bm{L}$ as the intensity of an ideal video sequence in the period of $[t_{s},t_{e}]$ containing continuous GS frames. A RS frame $I_{r}$ can be regarded as a special composition of a series of GS frames. We denote the $t^{r}_{s}$ and $t^{r}_{e}$ as the start and end of the exposure time of RS frame $I_{r}$, respectively. The interval of exposure time between every two adjacent lines is $\Delta t=(t^{r}_{s}-t^{r}_{e})/(H-1)$, where $H$ is the height of video frames. Therefore, the exposure time of row $h$ of the RS frame can be recorded as $t_{h}=t^{r}_{s}+h\times\Delta t$. Then the raw $h$ of RS frame $I_{rs}$ can be formulated as Eq. 1, where $I_{t_{h}}[h]$ refers to the row $h$ of the GS frame captured at $t_{h}$.
@@
IV Proposed Framework
Overview: The overall framework of the proposed approach is depicted in Fig. 2, which can be divided into three parts: (I) Displacement Field Estimation, (II) Latent GS frames Generation, and (III) Reciprocal Reconstruction. As the oversimplified linear motion model fails in complex non-linear motions, we first introduce DF which contains non-linear motion during the exposure time and bridge the gap between RS frames and GS frames (Sec. IV-A), followed by describing how to generate latent GS frames by RS frames and DF (Sec. IV-B). Since the mapping from RS to GS is highly under-constrained, we couple it with the inverse mapping (GS-to-RS) and RS frame warping (RS-to-RS) for self-supervision (Sec. IV-C). The inputs of our framework are two consecutive RS frames($I_{r_{0}}$, $I_{r_{1}}$) and their corresponding events, while the outputs are continuous latent global shutter (GS) frames at arbitrary frame rate, as shown in Fig. 3.
IV-A Displacement Field Estimation (DFE)
@@
To address this issue, we leverage the high-temporal resolution of events to estimate the optical flow within $[t_{0},t_{1}]$. We divide the events into $T+1$ time bins, and each time bin is further divided into $N$ voxel grids [53, 63]. Based on this, we can get the event representation with a dimension of $(T+1)\times N\times H\times W$. We denote the event-based optical flow estimation function as $f_{flow}$, and the dimension of the estimated optical flow set is $2\times T\times H\times W$. In practice, we estimate a set of optical flows as the $\mathbf{D}$ by E-RAFT  [55], as it has shown promising results in handling challenging motion scenes. In this way, we can fit the nonlinear motion with the linear motions of $T$ segments. If we know the location of a pixel at time $t_{0}$ and there is a considerable time gap between $t_{0}$ and $t_{n}$, we can obtain the location of this point at time $t_{n}$ by adding up its displacement in DF over the time interval, as Eq. 2.
@@
$\mathbf{L}_{field}=\frac{1}{T}\sum_{i=1}^{T}\left((\nabla_{x}\mathbf{D}[i])^{2}+(\nabla_{y}\mathbf{D}[i])^{2}\right)$ (3)
@@
Figure 4: An illustration of weight maps of RS-to-GS and RS-to-RS for the displacement field. Blue represents a positive value, indicating forward movement. Red represents a negative value, indicating backward movement. Each item in the weight map uses uniform sample $S_{h}\times S_{t}$ points to estimate its weight. Two time points define the rolling shutter exposure: the start time $t_{rs}$ and the end time $t_{re}$. In contrast, the global shutter frame is exposed at a single time point $t_{g}$. Subfigures (a), (b), and (c) provide detailed examples: (a) shows the transformation from an RS frame to a GS frame, (b) shows an RS-to-RS mapping, and (c) illustrates the mapping from an RS frame to a preceding GS frame.
@@
Weight map Design for RS GS transformation We designed the weight map based on the exposure mechanism of the RS and GS to describe the transition between any two frames. The exposure model of the rolling shutter frame and events are shown in Fig.4. The red solid line represents the start time of the frame exposure, and the red dashed line represents the end time of the frame exposure. It can be easily found that the exposure time is much shorter than the rolling shutter time. Long exposure time can lead to blurring, which complicates the research question of frame interpolation, and we do not consider the long exposure in the paper. Therefore, we approximate the rolling shutter frame and the global shutter frame as planes. Obviously, the plane of the global shutter can be described as $t=t_{g}$, where $t_{g}$ is the global shutter exposure timestamp. The plane of the rolling shutter is parallel to the W-axis, row of the image, and passes through two points $(h,t_{rs})$ and $(0,t_{re})$ at the same time, where $h$ is the height of frames, and $t_{rs}$ and $t_{re}$ are begin time and end time of rolling shutter frame exposure. Given two planes, we use uniform sampling to calculate the weights of each bin in each weight map. For each sampling point, we calculate whether it is on the left or right of the plane by computational geometry. Fig. 4 shows the projection of more weight maps in the $H\times T$ dimension. Fig. 6 shows the outputs of our framework and the visualization of warp weight in an aerial scene.
@@
Figure 5: The illustration of latent GS generation from RS frames $I_{r_{0}}$, $I_{r_{1}}$ and displacement field $\mathbf{D}$. The dash boxes indicate the output.
Figure 6: Evaluation of our framework on the UAV-RS dataset, we utilize Fastec simulation dataset [36] for training and test on this dataset. Displayed herein are the outcomes of both rolling shutter correction and frame interpolation. Additionally, a visualization of the warp weight is presented. From the visualization, it’s evident that our framework yields generalization performance on the UAV dataset with deformation correction and frame interpolation. For more vivid depict of this scene, please visit the accompanying video in the supplementary material.
IV-B Latent Global Shutter Frame Generation